# A simplified proof of the CSP Dichotomy Conjecture and XY-symmetric operations

### v1

@@ 1 Introduction
Finally, the new theory connected the two main ideas of Zhuk’s proof: the idea of strong subalgebras and the idea of bridges and connectedness. Originally, they lived separate lifes. Using strong subalgebras and reductions we showed that all the relations have the parallelogram property, which gives us an irreducible congruence for every constraint and its variable. Then manipulating with the instance we tried to connect the congruences (variables) with bridges. In the new theory, bridges appear naturally from strong/linear subuniverses: whenever a restriction to strong/linear subuniverses gives an empty set, it immediately gives a bridge between congruences defining these subuniverses.
Using this new theory we obtain two results presented in the next two subsections.
@@ 1.1 A simplified proof of the CSP Dichotomy Conjecture
Suppose $\Gamma$ is a constraint language preserved by a WNU operation $w$. Then each $D_{x}$ of size at least 2 has a strong subset (subuniverse) or an equivalence relation $\sigma$ such that $D_{x}/\sigma\cong\mathbb{Z}_{p}$ for some $p$.
@@
      using recursion calculate $\phi^{-1}(\mathcal{I}_{i})$ for every $i$;
@@
      $\phi^{-1}(\mathcal{I})=\phi^{-1}(\mathcal{I}_{1})\cup\dots\cup\phi^{-1}(\mathcal{I}_{r})$.
@@
      the equation defining $\varphi^{-1}(\mathcal{I})$ is $\sum_{i\in J}(y_{i}-a_{i})/(b_{i}-a_{i})=1$.
@@
6:         $\triangleright$ remove unnecessary constraint from $\mathcal{I}^{\prime}$
7:              for $C\in\mathcal{I}^{\prime}$ do
8:                  if $\varphi^{-1}(\mathcal{I}^{\prime}\setminus\{C\})\neq\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}$ then $\triangleright$ using (p1)
9:                       :=$\mathcal{I}^{\prime}:=\mathcal{I}^{\prime}\setminus\{C\}$
10:         until nothing changed
11:         $F:=\varphi^{-1}(\mathcal{I}^{\prime})$triangleright using (p2) or (p3)
12:         if $F=\varnothing$ then return “No”
13:         $m:=$ the dimension of $F$
14:         $\psi:=$ a bijective linear map $\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}\to F$$\triangleright$ $p_{1},\dots,p_{m}$ are also updated
15:         $\varphi:=\varphi\circ\psi$
16:     return “Yes”
The function SolveLinear solving the remaining case works as follows (see the pseudocode). We start with a bijective linear map $\varphi:\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}\to D_{x_{1}}/\sigma_{x_{1}}\times\dots\times D_{x_{n}}/\sigma_{x_{n}}$. We gradually reduce the dimension $m$ maintaining the property that $\operatorname{Im}{\varphi}$ contains all the solutions of $\mathcal{I}$. We stop when $\varphi^{-1}(\mathcal{I})=\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}$ or $\varphi^{-1}(\mathcal{I})$ is empty. First, we make a copy $\mathcal{I}^{\prime}$ of the instance $\mathcal{I}$ and remove all the constraints from $\mathcal{I}^{\prime}$ that can be removed so that $\mathcal{I}^{\prime}$ maintains the property $\varphi^{-1}(\mathcal{I}^{\prime})\neq\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}$. This property can be checked in polynomial time using (p1). If we cannot remove any other constraint, by Informal Claim 3 we have one of the following cases: either $\mathcal{I}^{\prime}$ is not linked, and we can calculate $\varphi^{-1}(\mathcal{I}^{\prime})$ using (p2); or $\mathcal{I}^{\prime}$ satisfies all the conditions of Informal Claim 3, and $\varphi^{-1}(\mathcal{I}^{\prime})$ has dimension $m-1$ or is empty. In the second case we can calculate $\varphi^{-1}(\mathcal{I}^{\prime})$ using (p3). Since $\mathcal{I}^{\prime}$ was obtained from $\mathcal{I}$ by removing some constraints, we have $\varphi^{-1}(\mathcal{I})\subseteq\varphi^{-1}(\mathcal{I}^{\prime})$. Thus, we found a smaller affine subspace $\varphi^{-1}(\mathcal{I}^{\prime})$ that still covers all the solutions of $\mathcal{I}$. It remains to replace $m$ with the dimension of $\varphi^{-1}(\mathcal{I}^{\prime})$ and update the linear map $\varphi$. Since we cannot reduce the dimension $m$ forever, we will eventually stop in one of the two cases: $\varphi^{-1}(\mathcal{I})=\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}$ or $\varphi^{-1}(\mathcal{I})$ is empty. First of them implies the existence of a solution for $\mathcal{I}$, the second implies that no solutions exist.
1.2 Existence of an XY-symmetric operations
@@
An $n$-ary operation $f$ is symmetric on a tuple of variables $(x_{i_{1}},\dots,x_{i_{n}})$ if it satisfies the identity $f(x_{i_{1}},\dots,x_{i_{n}})=f(x_{i_{\sigma(1)}},\dots,x_{i_{\sigma(n)}})$ for every permutation $\sigma$ on $\{1,2,\dots,n\}$. For instance, an operation $f$ is symmetric on $(x,\dots,x,y)$ if and only if $f$ is a WNU operation.
An operation is called XY-symmetric if it is symmetric on $(\underbrace{x,\dots,x}_{i},y,\dots,y)$ for any $i$.
As it follows from the definition, an XY-symmetric operation satisfies much more identities than a WNU operation. Nevertheless, we managed to prove that an XY-symmetric operation can always be derived from a WNU operation. To formulate the precise statement we will need an extra notation. A set of operations is called a clone if it is closed under composition and contains all projections. For a set of operations $M$ by $\operatorname{Clo}(M)$ we denote the clone generated by $M$.
@@
$f(x_{1},x_{2},x_{3})=\begin{cases}x_{1}+x_{2}+x_{3}\;(\mathrm{mod}\;\;2)&\text{if $x_{1},x_{2},x_{3}\in\{0,1\}$}\\ 2&\text{if $x_{1}=x_{2}=x_{3}=2$}\\ \text{first non-2 in $x_{1},x_{2},x_{3}$}&\text{otherwise}\end{cases}$
The following lemma shows that even if we try to generalize XY-symmetric operations to some tuples with $x$, $y$, and $z$ we fail.
@@
The operation $f$ is conservative, and behaves as linear sum on $\{0,1\}$, as $\mathrm{min}$ on $\{0,2\}$ and $\{1,2\}$. Therefore composing $f$ we can get a WNU operation of any add arity.
Also, $f$ has the following properties. Whenever we substitute 00 or $1$ in $f$ we obtain either 00, or $1$. Whenever we substitute 2 into some arguments, we get an operation whose restriction to $\{0,1\}$ is a linear operation. Finally, the operation preserves the sets $\{0,2\}$ and $\{1,2\}$. These three properties imply that if we put 2 to some arguments of a term operation $g\in\operatorname{Clo}(\{f\})$ and restrict the obtained operation to $\{0,1\}$ we get an idempotent linear operation on $\{0,1\}$.
@@
Notice that both Zhuk’s and Bulatov’s algorithms for the CSP are not universal in the sense that the algorithms work only if the domain is fixed and, therefore, all the algebraic properties are known. It would be great to find a universal algorithm for all tractable CSPs. Recently the importance of symmetric operations was rediscovered while studying the limits of universal algorithms for the CSP and its variation, called the Promise CSP [16, 2, 11]. For instance, the algorithm known as BLP+AIP solves $\operatorname{CSP}(\Gamma)$ if and only if for $\Gamma$ has infinitely many symmetric polymorphisms [11].
We believe that Theorem 2 can be further generalized and finally we will get enough symmetric operations to make some universal algorithm work. Theorem 2 already gives us some implications that can be viewed as a tiny step in this direction:
@@
  • if $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ is solved by reducing to $\operatorname{CSP}(\mathbb{C})$, where $\mathbb{A}\to\mathbb{C}\to\mathbb{B}$, $\mathbb{C}$ is finite, and $|A|=2$, then $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ is solvable by BLP+AIP (see [23, 2] for more information about Promise CSP).
@@ 3 Proof of the CSP Dichotomy Conjecture
In this section we prove Theorems 42 and 43 that show the correctness of Zhuk’s algorithm for the CSP. We start with Subsection 3.1, where we give additional definitions such as irreducible, linked, and crucial instances of the CSP. Crucial instances are the instances that have no solutions but any weakening of the instance (like removing a constraint) gives an instance with a solution. We will show any constraint in a crucial and consistent enough instance has the parallelogram property. This allows us to define a congruence for every constraint and its variable and talk about connectedness of the variables by bridges. Also, we explain how we weaken the instance: usually we just replace a constraint by a weaker constraint but sometimes we also need to disconnect two constraints by adding an additional variable, which leads us to the notion of Expanded Coverings.
@@ 3.2 Auxiliary statements
Define a new reduction $D^{(\bot)}$ by $D_{y}^{(\bot)}=G$ and $D_{x}^{(\bot)}=D_{x}^{(\top)}$ for every $x\neq y$. We extend the reduction $D^{(\top)}$ and $D^{(\bot)}$ so that the reductions on $x_{i}$ and $x_{i}^{\prime}$ coincide and $D_{x_{0}^{\prime\prime}}^{(\top)}=D_{x_{0}^{\prime\prime}}^{(\bot)}=D_{x_{0}}$. Since the instance $\Theta\wedge\mathcal{I}\wedge\omega(x_{0},x_{0}^{\prime\prime})$ has a subdirect solution set, by Corollary 21 the types $\mathcal{T}$ and $\mathcal{T}_{0}$ are the same. Moreover, if $\mathcal{T}\in\{\mathcal{L},\mathcal{PC}\}$, there exists a bridge $\delta^{\prime}$ from $\xi$ to $\nu$ such that $\widetilde{\delta}^{\prime}\supseteq\Theta(z^{\prime},x_{0}^{\prime\prime})\circ\mathcal{I}(x_{0},y)\supseteq\mathcal{I}(z,y)$ (if $z=y$ then it is just a trivial reflexive bridge). Let $F$ be set of possible values of $x_{0}^{\prime\prime}$ in the solutions of $\Theta^{(2)}\wedge\mathcal{I}^{(\top)}\wedge\omega(x_{0},x_{0}^{\prime\prime})$. In other words, $F=\Theta^{(2)}(x_{0}^{\prime\prime})\cap(\mathcal{I}^{(\top)}(x_{0})\circ\omega)$. Since the variable $x_{0}^{\prime\prime}$ only appears in the constraint $C$ and $\omega^{*}$ we have $F\circ\omega=F$. By Corollaries 17(r1) $F\lll D_{x_{1}}$. By Lemma 12 we find a single block $E$ of $\omega$ such that $\{E\}\lll^{D_{x_{0}}/\omega}F/\omega$. By Corollary 16(f) $E\lll F$.

### v2

@@ 1 Introduction
Finally, the new theory connected the two main ideas of Zhuk’s proof: the idea of strong subalgebras and the idea of bridges and connectedness. Originally, they lived separate lifes. Using strong subalgebras and reductions we showed that all the relations have the parallelogram property, which gives us an irreducible congruence for every constraint and its variable. Then manipulating with the instance we tried to connect the congruences (variables) with bridges. In the new theory, bridges appear naturally from strong/linear subuniverses: whenever a restriction to strong/linear subuniverses gives an empty set, it immediately gives a bridge between congruences (equivalence relations) defining these subuniverses.
Using the new theory we obtain two results presented in the next two subsections.
@@ 1.1 A simplified proof of the CSP Dichotomy Conjecture
Suppose $\Gamma$ is a constraint language preserved by a WNU operation $w$. Then each $D_{x}$ of size at least 2 has a strong subset (subuniverse) or an equivalence relation $\sigma$ such that $D_{x}/\sigma\cong\mathbb{Z}_{p}$ for some prime $p$.
@@
      using recursion calculate $\varphi^{-1}(\mathcal{I}_{i})$ for every $i$;
@@
      $\varphi^{-1}(\mathcal{I})=\varphi^{-1}(\mathcal{I}_{1})\cup\dots\cup\varphi^{-1}(\mathcal{I}_{r})$.
@@
      the equation defining $\varphi^{-1}(\mathcal{I})$ is $\sum_{i\in J}(y_{i}-a_{i})/(b_{i}-a_{i})=1$.  [^1] The fact that different variables $y_{i}$ take on values from different fields $\mathbb{Z}_{p_{i}}$ is not a problem as $J$ may contain only variables on the same field.
@@
6:         for $C\in\mathcal{I}^{\prime}$ do $\triangleright$ remove unnecessary constraint from $\mathcal{I}^{\prime}$
7:             if $\varphi^{-1}(\mathcal{I}^{\prime}\setminus\{C\})\neq\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}$ then $\triangleright$ using (p1)
8:                 :=$\mathcal{I}^{\prime}:=\mathcal{I}^{\prime}\setminus\{C\}$
9:         $F:=\varphi^{-1}(\mathcal{I}^{\prime})$triangleright using (p2) or (p3)
10:         if $F=\varnothing$ then return “No”
11:         $m:=$ the dimension of $F$
12:         $\psi:=$ a bijective linear map $\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}\to F$$\triangleright$ $p_{1},\dots,p_{m}$ are also updated
13:         $\varphi:=\varphi\circ\psi$
14:    return “Yes”
The function SolveLinear solving the remaining case works as follows (see the pseudocode). We start with a bijective linear map $\varphi:\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}\to D_{x_{1}}/\sigma_{x_{1}}\times\dots\times D_{x_{n}}/\sigma_{x_{n}}$. We gradually reduce the dimension $m$ maintaining the property that $\operatorname{Im}{\varphi}$ contains all the solutions of $\mathcal{I}$. We stop when $\varphi^{-1}(\mathcal{I})=\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}$ or $\varphi^{-1}(\mathcal{I})$ is empty. First, we make a copy $\mathcal{I}^{\prime}$ of the instance $\mathcal{I}$ and remove all the constraints from $\mathcal{I}^{\prime}$ that can be removed so that $\mathcal{I}^{\prime}$ maintains the property $\varphi^{-1}(\mathcal{I}^{\prime})\neq\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}$. This property can be checked in polynomial time using (p1). If we cannot remove any other constraint, by Informal Claim 3 we have one of the following cases: either $\mathcal{I}^{\prime}$ is not linked and we can calculate $\varphi^{-1}(\mathcal{I}^{\prime})$ using (p2); or $\mathcal{I}^{\prime}$ satisfies all the conditions of Informal Claim 3, and $\varphi^{-1}(\mathcal{I}^{\prime})$ has dimension $m-1$ or is empty. In the second case we can calculate $\varphi^{-1}(\mathcal{I}^{\prime})$ using (p3). Since $\mathcal{I}^{\prime}$ was obtained from $\mathcal{I}$ by removing some constraints, we have $\varphi^{-1}(\mathcal{I})\subseteq\varphi^{-1}(\mathcal{I}^{\prime})$. Thus, we found a smaller affine subspace $\varphi^{-1}(\mathcal{I}^{\prime})$ that still covers all the solutions of $\mathcal{I}$. It remains to replace $m$ with the dimension of $\varphi^{-1}(\mathcal{I}^{\prime})$ and update the linear map $\varphi$. Since we cannot reduce the dimension $m$ forever, we will eventually stop in one of the two cases: $\varphi^{-1}(\mathcal{I})=\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{m}}$ or $\varphi^{-1}(\mathcal{I})$ is empty. First of them implies the existence of a solution for $\mathcal{I}$, the second implies that no solutions exist.
1.2 Existence of XY-symmetric operations
@@
An $n$-ary operation $f$ is called symmetric on a tuple of variables $(x_{i_{1}},\dots,x_{i_{n}})$ if it satisfies the identity $f(x_{i_{1}},\dots,x_{i_{n}})=f(x_{i_{\sigma(1)}},\dots,x_{i_{\sigma(n)}})$ for every permutation $\sigma$ on $\{1,2,\dots,n\}$. For instance, an operation $f$ is symmetric on $(x,\dots,x,y)$ if and only if $f$ is a WNU operation. An operation is called XY-symmetric if it is symmetric on $(\underbrace{x,\dots,x}_{i},y,\dots,y)$ for any $i$. An operation $f$ is called idempotent if $f(x,x,\dots,x)=x$.
As it follows from the definition, an XY-symmetric operation satisfies much more identities than a WNU operation. Nevertheless, we managed to prove that an XY-symmetric operation can always be derived from a WNU operation. To formulate the precise statement we will need a definition of a clone. A set of operations is called a clone if it is closed under composition and contains all projections. For a set of operations $M$ by $\operatorname{Clo}(M)$ we denote the clone generated by $M$.
@@
Suppose $f$ is a WNU operation of an odd arity $n$ on a finite set. Then there exists an XY-symmetric operation $f^{\prime}\in\operatorname{Clo}(\{f\})$ of arity $n$.
Theorem 2 extends known characterization of finite Taylor algebras.
Corollary 3.
Suppose $\mathbf{A}$ is a finite idempotent algebra. Then the following conditions are equivalent:
  1. 1.
    $\mathbf{A}$ is a Taylor algebra (satisfies nontrivial identities);
  2. 2.
    there does not exist an algebra $\mathbf{B}\in\mathrm{HS}(\mathbf{A})$ of size 2 whose operations are projections [13];
  3. 3.
    $\mathbf{A}$ has a WNU term operation of any prime arity $p>|A|$ [29];
  4. 4.
    $\mathbf{A}$ has a cyclic term operation of any prime arity $p>|A|$, i.e. an operation $c_{p}$ satisfying
[…]
  5. 5.
    $\mathbf{A}$ has a Siggers term operation, i.e an operation $f$ satisfying
[…]
  6. 6.
    $\mathbf{A}$ has an XY-symmetric term operation of any prime arity $p>|A|$.
Composing a cyclic operation $c_{n}$ of arity $n$ and an XY-symmetric operation $f$ of arity $n$ we can get an operation which is simultaneously cyclic and XY-symmetric:
$f^{\prime}(x_{1},\dots,x_{n}):=c_{n}(f(x_{1},\dots,x_{n}),f(x_{2},\dots,x_{n},x_{1}),\dots,f(x_{n},x_{1},\dots,x_{n})).$
Hence, conditions 4 and 6 of Corollary 3 give an infinite sequence of cyclic XY-symmetric operations, which are the most symmetric operations known to be in every finite Taylor algebra.
@@
$f(x_{1},x_{2},x_{3})=\begin{cases}x_{1}+x_{2}+x_{3}\;(\mathrm{mod}\;\;2),&\text{if $x_{1},x_{2},x_{3}\in\{0,1\}$}\\ 2,&\text{if $x_{1}=x_{2}=x_{3}=2$}\\ \text{first element different from 2 in $x_{1},x_{2},x_{3}$},&\text{otherwise}\end{cases}$
Note that the clone generated by $f$ is a minimal Taylor clone and its operations were completely described in [22]. The following lemma shows that even if we try to generalize XY-symmetric operations to some tuples with $x$, $y$, and $z$ we fail.
@@
The operation $f$ is conservative (always returns one of the coordinates), and behaves as a linear sum modulo 2 on $\{0,1\}$, as $\mathrm{min}$ on $\{0,2\}$ and $\{1,2\}$. Let us show how to derive a WNU (and even XY-symmetric) operation of any odd arity. Put
$f_{3}:=f,\;\;\;\;\;\;\;\;\;f_{2n+1}(x_{1},\dots,x_{2n+1}):=f_{2n-1}(f(x_{1},x_{2},x_{3}),x_{4},\dots,x_{2n+1}).$
It follows immediately from the definition that $f_{2n+1}$ is symmetric on $\{0,1\}$, on $\{0,2\}$, and on $\{1,2\}$.
Also, $f$ has the following properties. Whenever we substitute 00 or $1$ in $f$ we obtain either 00, or $1$. Whenever we substitute 2 into some arguments, we get an operation whose restriction to $\{0,1\}$ is a linear operation (in fact a projection). Finally, the operation preserves the sets $\{0,2\}$ and $\{1,2\}$. These three properties imply that if we put 2 to some arguments of a term operation $g\in\operatorname{Clo}(\{f\})$ and restrict the obtained operation to $\{0,1\}$ we get an idempotent linear operation on $\{0,1\}$.
@@
Notice that both Zhuk’s and Bulatov’s algorithms for the CSP are not universal in the sense that the algorithms work only if the domain is fixed and, therefore, all the algebraic properties are known. It would be great to find a universal algorithm for all tractable CSPs. Recently the importance of symmetric operations was rediscovered while studying the limits of universal algorithms for the CSP and its variation, called the Promise CSP [17, 2, 12]. For instance, the algorithm known as BLP+AIP solves $\operatorname{CSP}(\Gamma)$ if and only if $\Gamma$ has infinitely many symmetric polymorphisms [12].
We believe that Theorem 2 can be further generalized, and finally we will get enough symmetric operations to make some universal algorithm work. Theorem 2 already gives us some implications that can be viewed as a tiny step in this direction:
@@
  • if $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ is solvable by reducing to a tractable $\operatorname{CSP}(\mathbb{C})$, where $\mathbb{A}\to\mathbb{C}\to\mathbb{B}$, $\mathbb{C}$ is finite, and $|A|=2$, then $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ is solvable by BLP+AIP (see [26, 2] for more information about Promise CSP).
To show the second claim we apply Theorem 2 to $\operatorname{Pol}(\mathbb{C})$ and obtain infinitely many XY-symmetric operations on $C$. Composing them with the homomorphisms $\mathbb{A}\to\mathbb{C}$ and $\mathbb{C}\to\mathbb{B}$ we obtain infinitely many symmetric polymorphisms $\mathbb{A}\to\mathbb{B}$, which implies that BLP+AIP solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ [12].
@@
1.3 History and acknowledgements
The first symmetric operations (WNU) that exist in every Taylor algebra appeared in [29], and the idea was to show that every symmetric invariant relation has a constant tuple. In [34] I showed the existence of a constant tuple in a symmetric relation gradually reducing the domain to strong subalgebras and keeping the property that the relation is symmetric. It turned out that the only reason why a constant tuple does not exist in a symmetric relation is a linear essence inside, for instance the relation $x_{1}+\dots+x_{p}=1$ does not have a constant tuple in $\mathbb{Z}_{p}$. Thus, the existence of a WNU term operation of an arity $n$ is reduced to a pure linear algebra question: does every affine symmetric subspace of $\mathbb{Z}_{p}^{n}$ have a constant tuple. Similarly, we could try to show the existence of a 2-WNU operation (symmetric on $(x,x,y,y,\dots,y)$). For the proof to work we need to show that every relation of arity (${n\choose 2}$ with symmetries coming from $n$-permutations has a constant tuple. This question is again reduced to a pure linear algebra question: does every affine (weak) symmetric subspace of $\mathbb{Z}_{p}^{n\choose 2}$ have a constant tuple. We worked on it together with Libor Barto, Michael Pinsker, and their students Johanna Brunar and Martin Boroš. Martin Boroš in his master thesis [9] proved that a constant tuple always exists if and only if $n\cdot{n\choose 2}$ is co-prime with $p$. Unfortunately this beautiful approach did not lead to a proof of existence of 2-WNU or XY-symmetric operations as it turned out that an XY-symmetric operation may exist even if the condition on the arity is not satisfied. Nevertheless, I am very thankful to Libor Barto and Michael Pinsker for the exciting play with a beautiful linear algebra and the ideas I took from this play.
I would also like to thank Stanislav Živný, Lorenzo Ciardo, and Tamio-Vesa Nakajima for very fruitful discussions about algorithms for the CSP based on linear programming and their limits. My understanding of what operations we need for the algorithms to work came to me during my visit of Oxford University. The impotence of symmetric operations for these algorithms motivated me to finish my research on XY-symmetric operations.
@@ 3 Proof of the CSP Dichotomy Conjecture
In this section we prove Theorems 43 and 44 that show the correctness of Zhuk’s algorithm for the CSP. We start with Subsection 3.1, where we give additional definitions such as irreducible, linked, and crucial instances of the CSP. Crucial instances are the instances that have no solutions but any weakening of the instance (like removing a constraint) gives an instance with a solution. We will show any constraint in a crucial and consistent enough instance has the parallelogram property. This allows us to define a congruence for every constraint and its variable and talk about connectedness of the variables by bridges. Also, we explain how we weaken the instance: usually we just replace a constraint by a weaker constraint but sometimes we also need to disconnect two constraints by adding an additional variable, which leads us to the notion of Expanded Coverings.
@@ 3.2 Auxiliary statements
Define a new reduction $D^{(\bot)}$ by $D_{y}^{(\bot)}=G$ and $D_{x}^{(\bot)}=D_{x}^{(\top)}$ for every $x\neq y$. We extend the reduction $D^{(\top)}$ and $D^{(\bot)}$ so that the reductions on $x_{i}$ and $x_{i}^{\prime}$ coincide and $D_{x_{0}^{\prime\prime}}^{(\top)}=D_{x_{0}^{\prime\prime}}^{(\bot)}=D_{x_{0}}$. Since the instance $\Theta\wedge\mathcal{I}\wedge\omega(x_{0},x_{0}^{\prime\prime})$ has a subdirect solution set, by Corollary 22 the types $\mathcal{T}$ and $\mathcal{T}_{0}$ are the same. Moreover, if $\mathcal{T}\in\{\mathcal{L},\mathcal{PC}\}$, there exists a bridge $\delta^{\prime}$ from $\xi$ to $\nu$ such that $\widetilde{\delta}^{\prime}\supseteq\Theta(z^{\prime},x_{0}^{\prime\prime})\circ\mathcal{I}(x_{0},y)\supseteq\mathcal{I}(z,y)$ (if $z=y$ then it is just a trivial reflexive bridge). Let $F$ be set of possible values of $x_{0}^{\prime\prime}$ in the solutions of $\Theta^{(2)}\wedge\mathcal{I}^{(\top)}\wedge\omega(x_{0},x_{0}^{\prime\prime})$. In other words, $F=\Theta^{(2)}(x_{0}^{\prime\prime})\cap(\mathcal{I}^{(\top)}(x_{0})\circ\omega)$. Since the variable $x_{0}^{\prime\prime}$ only appears in the constraint $C$ and $\omega^{*}$ we have $F\circ\omega=F$. By Corollaries 18(r1) $F\lll D_{x_{1}}$. By Lemma 13 we find a single block $E$ of $\omega$ such that $\{E\}\lll^{D_{x_{0}}/\omega}F/\omega$. By Corollary 17(f) $E\lll F$.